OpenStudy (anonymous):

What is the slope of the line passing through the point (3, −4) and perpendicular to the line y = 3x + 5? Write your answer as a fraction.

OpenStudy (3psilon):

So the line is perpendicular to y = 3x + 5 that means the slope of the new line is the inverse of that slope. So the slope of the new like is -1/3(x)

OpenStudy (3psilon):

Line*

OpenStudy (anonymous):

y=-1/3x-3 and here is the equation for the line

OpenStudy (3psilon):

Use y = mx + b to solve for the y-intercept (b) Just plug in the values (x,y) and the new slope m.

OpenStudy (3psilon):

Then you'll be able to write it in y = mx +b form